[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Tue Aug 08, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Tue Aug 8 19:30:06 UTC 2017


HSU Shift Log for Tue Aug 08 2017 15:30:00 GMT-0400 (EDT)

Shift: H

Forecaster: J Cangialosi

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues: Requested GOES-16 meso for the following times: 00-15  
Z Aug 9 15 Z Aug 9 - 13 Z Aug 10. Not sure if either of these requests will  
be granted.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170808/39bc22da/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list